How many ways can 12 people out of 300 be selected for a jury?
Found 2 solutions by checkley75, Edwin McCravy:
Answer by checkley75(3666)

300*299*298*297*296*295*294*293*292*291*290*289=4.25e29

Answer by Edwin McCravy(20054)


How many ways can 12 people out of 300 be selected for a jury?

This is the combinations of 300 things taken 12 at a time.

This is sometimes written C(300,12) and sometimes 300C12

Checkley's answer is wrong as he thought it was permutations,
but because the order of the members of a jury is not important,
we must use combinations, not permutations, which means we 
must divide his answer by 12! or 12·11·10·9·8·7·6·5·4·3·2·1 =
479001600.

So Checkley should have had

 300*299*298*297*296*295*294*293*292*291*290*289     4.252707522E29
------------------------------------------------- = ----------------- =
          12*11*10*9*8*7*6*5*4*3*2*1                    479001600


 4.252707522 × 1029
--------------------- = 8.878274148 × 1020 
       479001600

Here is how to do it on a TI-83 or TI-84 calculator:

Type in 300
Press MATH
Press | 
Press 3       (You should see this on the main screen: 300 nCr
Type in 12      (You should see this on the main screen: 300 nCr 12
Press ENTER

Read this:  8.878274148E20

This means 8.878274148 × 1020

or about 

887,827,414,800,000,000,000

887 quintillion, 827 quadrillion, 414 trillion, 800 billion

Edwin